Starting from rest, a 12-cm-diameter compact disk takes 2.9s to reach its operating angular velocity of 2500rpm . Assume that the angular acceleration is constant. The disk's moment of inertia is 2.5
Answer #1

w0=0:
r=0.06 m;t=2.9 sec;w=2500 =2500*6.28/3600=4.36 rad./s
I=2.5*10^5 kg*m^2

first we hsould have to calculate angular acceleration (alpha)?

w=w0+alpha*t
alpha=w/t=4.36/2.9=1.5038 rad/s^2


A)torque=I*alpha=2.5*10^5*1.5038=3.759*10^5 N.m

B)w^2-wo^2=2*alpha*theta
4.36^2=2*1.5038*theta
theta =6.320 rad